# VideoLLaMA 2 Advancing Spatial-Temporal Modeling and Audio Understanding in Video-LLMs

### v1

@@ Abstract
In this paper, we present the VideoLLaMA 2, a set of Video Large Language Models (Video-LLMs) designed to enhance spatial-temporal modeling and audio understanding in video and audio-oriented tasks. Building upon its predecessor, VideoLLaMA 2 incorporates a tailor-made Spatial-Temporal Convolution (STC) connector, which effectively captures the intricate spatial and temporal dynamics of video data. Additionally, we integrate an Audio Branch into the model through joint training, thereby enriching the multimodal understanding capabilities of the model by seamlessly incorporating audio cues. Comprehensive evaluations on multiple-choice video question answering (MC-VQA), open-ended video question answering (OE-VQA), and video captioning (VC) tasks demonstrate that VideoLLaMA 2 consistently achieves competitive results among open-source models and even gets close to some proprietary models on several benchmarks. Furthermore, VideoLLaMA 2 exhibits reasonable improvements in audio-only and audio-video question-answering (AQA & OE-AVQA) benchmarks over existing models. These advancements underline VideoLLaMA 2’s superior performance in multimodal comprehension, setting a new standard for intelligent video analysis systems. All models are public to facilitate further research.
@@ 1 Introduction
In recent years, the field of artificial intelligence (AI) has achieved significant advancements (OpenAI, 2023a; Google, 2023; Anthropic, 2024), profoundly transforming industries and societal functions across the board. Models capable of image recognition (Bai et al., 2023; Dong et al., 2024; Chen et al., 2023d; Liu et al., 2024c; Chen et al., 2023b; Lin et al., 2023c; Young et al., 2024) and photorealistic image generation (Esser et al., 2024; Saharia et al., 2022; Ramesh et al., 2021) have approached near-human capabilities, catalyzing major breakthroughs in sectors such as medical imaging (Li et al., 2024; Tu et al., 2024) and autonomous driving (Xu et al., 2023; Jin et al., 2023a). Despite these successes, the domain of video understanding and generation (Kondratyuk et al., 2023; Menapace et al., 2024; Kondratyuk et al., 2023; Brooks et al., 2024) remains relatively nascent. Unlike static images, videos incorporate temporal dynamics and synchronous audio streams, significantly enriching the information content. This integration of continuous audio-visual data complicates extracting and interpreting meaningful patterns, amplifying data complexity and introducing unique computational challenges.
While Image Large Language Models (Image-LLMs) (Alayrac et al., 2022; Li et al., 2023a; Zhu et al., 2023b; Liu et al., 2024c; Bai et al., 2023; Chen et al., 2023d; Dong et al., 2024) processing static images have matured with impressive capabilities, Video Large Language Models (Video-LLMs) (Li et al., 2023b; Zhang et al., 2023; Maaz et al., 2023a; Lin et al., 2023a; Wang et al., 2024b; Liu et al., 2024a) lag notably behind due to inherent complexities. The primary challenge in video understanding lies in the temporal dynamics—recognizing visual patterns while understanding changes over time and correlating these with synchronous audio inputs. These temporal dynamics complicate the accurate prediction of future states and the understanding of complex scenarios, such as interactions among multiple entities or subtle environmental changes. Moreover, the integration of audio with visual data, essential for comprehensive understanding, remains underdeveloped in current models, limiting their effectiveness.
@@ 2.1 Architecture
With an established foundation in the visual domain, our exploration extends into the auditory realm to enhance the multimodal capabilities of VideoLLaMA. Initially, audio signals undergo a preprocessing step where they are transformed into fbank spectrograms with 128 frequency bins. To effectively harness these preprocessed audio signals, we integrate BEATs (Chen et al., 2023c), a cutting-edge audio encoder, known for its exceptional ability to capture detailed audio features and temporal dynamics. These features are then processed through a simple linear layer to align with the dimension of LLMs, therefore providing a more cohesive understanding of the video content when combined with the visual and acoustic modalities. By incorporating BEATs into VideoLLaMA, we address the challenge of synchronizing audio-visual data points. The encoder’s ability to capture temporal dynamics aligns with the STC Connector employed in the visual branch, ensuring a seamless integration of audio-visual features.
@@
Like its predecessor, VideoLLaMA 2 adopts the instruction-following large language models (LLMs) as its language decoder. We do not extensively search the optimal LLMs for VideoLLaMA 2 but use Mistral-Instruct (Jiang et al., 2023) and Mixtral-Instruct (Jiang et al., 2024) across all of the experiments. We leave the exploration of other popular LLMs, such as Gemma-IT (Team et al., 2024), LLaMA3-Instruct111https://github.com/meta-llama/llama3 and Qwen2-Instruct (Qwen, 2024), for future work.
@@
Sptial-Temporal Convolution Connector
The overall architecture and the workflow of our STC connector are depicted in Figure 2. We primarily follow three principles when designing the video-language connector: 1) Maintaining the spatial-temporal order in the output visual tokens; 2) Reducing the number of spatial-temporal tokens; 3) Alleviating information loss during spatial-temporal downsampling. ❶ The first principle guides us to avoid using resampler architecture (Li et al., 2023a; Alayrac et al., 2022) because the resampling operation does not guarantee the preservation of spatial-temporal order. This may cause suboptimal convergence because autoregressive models (i.e., LLM backbone) are highly dependent on consistent token order between training and inference. Therefore, we only consider the operations of convolution or pooling when building our connector. ❷ According to the second principle, we insert 3D downsample operator to compress spatial-temporal tokens. ❸ To complement the information loss caused by the spatial-temporal downsampling, we insert RegStage (a strong convolution block (Radosavovic et al., 2020)) before and after spatial-temporal downsampling. To empirically investigate the effectiveness of the designs above, we establish a quick but reasonable architectural search using the training data from Video-LLaVA (Lin et al., 2023a). We select three representative video understanding benchmarks: Egoschema (Mangalam et al., 2024), MV-Bench (Li et al., 2023c) and ActivityNet-QA (Yu et al., 2019) as the testbed222See Section 4.1 for benchmark details and evaluation protocols. and the comparison results are listed in Table 1. As can be seen, 3D convolution together with the ReStage block (i.e., the green line), which forms our STC connector, works the best in terms of the average performance. Another interesting finding is that almost all of the 3D downsampling designs perform better than the 2D ones on Egoschema, suggesting that the early fusion of frame-level features is beneficial for long video understanding.
@@ 3 Training
| Video-Text | InternVid-10M (Wang et al., 2023) | 10M | 650K | 6.5% |
@@ 3.1.1 Pre-training
For the pre-training stage, we utilize a large-scale, weakly labeled, web-crawled dataset of image-text and video-text pairs, sourced from several publicly accessible databases. The video-text sources include Panda-70M (Chen et al., 2024b), VIDAL-10M (Zhu et al., 2023a), WebVid-10M (Bain et al., 2021), and InternVid-10M (Wang et al., 2023), while the image-text sources include CC-3M (Changpinyo et al., 2021) and DCI (Urbanek et al., 2023).
@@ 3.1.2 Multi-task Fine-tuning
In the second stage of multi-task fine-tuning, we incorporate high-quality, fine-grained multi-modal annotations using both video-text and image-text data. As summarized in Table 3 , VideoLLaMA 2 is fine-tuned on four tasks simultaneously. For video captioning, we use training samples from VideoChat (Li et al., 2023b) and in-house collected data. For video classification and VQA tasks, we utilize a mixture of publicly available datasets, including Kinetics-710 (Kay et al., 2017), SthSthv2 (Goyal et al., 2017), NExtQA (Xiao et al., 2021), CLEVRER (Yi et al., 2019), EgoQA (Fan, 2019), Tgif (Li et al., 2016), WebVidQA(Yang et al., 2021), RealworldQA (X.AI, 2024), and Hm3d (Ramakrishnan et al., 2021). To enhance instruction-following capabilities, we collect and integrate data from multiple frontier Video-LLMs, such as Vally (Luo et al., 2023), VideoChatGPT (Maaz et al., 2023b), VideoChat (Li et al., 2023b), VTimeLLM (Huang et al., 2023), and VideoChat2 (Li et al., 2023c) to improve generalization. Additionally, we mix image captioning data from ShareGPT4V (Chen et al., 2023b) and image VQA and instruction-following data from LLaVA (Liu et al., 2023b, a) to maintain the capabilities in understanding static visual concepts.
@@ 3.2.2 Multi-task Fine-tuning
The second stage of audio-language training aims to enhance the versatility and applicability of our model through multi-task learning, involving a variety of datasets tailored to different audio processing tasks. ClothoAQA (Lipping et al., 2022b), with about 1,500 entries, is utilized for refining our model’s capabilities in question answering based on audio cues, with each sample enriched by six associated questions sourced through crowdsourcing. The instruction tuning phase also capitalizes on the continued use of WavCaps (Mei et al., 2023), alongside AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020), which together contribute about 453,000 audio samples for audio-text training. MusicCaps (Agostinelli et al., 2023), comprising roughly 5,400 entries, extends our model’s exposure to the musical domain, specializing in music-based captioning. For the purpose of sound event classification, we engage with VGGSound (Chen et al., 2020), which offers over 190,000 labeled audio clips of diverse acoustic events. UrbanSound8k (Salamon et al., 2014) and ESC50 (Piczak, 2015) provide additional layers of complexity with 8000+ and 2,000 urban and miscellaneous environmental sounds, respectively. TUT2017 (Mesaros et al., 2017) and VocalSound (Gong et al., 2022) enrich our dataset collection with more than 4,000 and 15,000 samples respectively, focusing on general sound events and human vocal sounds classification, thus broadening our model’s acoustic perceptual skills.
@@ 3.3 Audio-Video Joint Trainng
The third phase shifts our focus towards the integration of audio and visual modalities, aiming to harness and understand the interactions between these two crucial aspects of multimodal perception. This phase incorporates audio-visual datasets including AVQA (Yang et al., 2022) and AVQA-music (Li et al., 2022b), featuring around 57,000 and 35,000 entries, respectively. These datasets are specifically designed for audio-visual question answering, challenging our model to not only perceive but also interpret cross-modal content effectively. The AVSD (Alamri et al., 2019) dataset, consisting of approximately 80,000 pairs, is pivotal for developing and refining audio-visual dialogue systems. Furthermore, VGGSound (Chen et al., 2020) reappears with a substantial subset dedicated to audio-visual classification tasks, reinforcing the dataset’s utility and versatility in our multimodal training framework. The integration of multimodal data continues with the inclusion of VideoInstruct100K, a rich dataset comprising about 100,000 video clips, each annotated with GPT-generated instruction-response pairs. This dataset is particularly designed to enhance the interactive capabilities of models by providing scenarios that require complex comprehension and response generation based on visual cues and verbal instructions. Complementing this, we incorporate the WebVid subset, which consists of about 230,000 video clips accompanied by concise, descriptive textual annotations. This dataset is specifically utilized to bolster the model’s understanding of visual-only instructions, offering a diverse array of contexts and scenarios.
@@
In this stage, audio tracks are extracted from videos and cut to align the video clips. These audio clips are then truncated or padded to the same duration as the audio-language traning stage. For videos that lack an audio track, we fill the waveform with zeros to ensure uniformity across all data samples. The video encoder and the audio encoder remain frozen while we optimize the audio/video projector, alongside the unfrozen LLM. The training objectives align with the instruction tuning stage, ensuring a coherent and effective progression in our multimodal training approach. By effectively leveraging the synchronized audio-visual data, VideoLLaMA 2 achieves a deeper understanding of multimodal content, thereby enhancing its performance across a spectrum of multimedia analysis tasks.

### v3

@@ Abstract
In this paper, we present VideoLLaMA 2, a set of Video Large Language Models (Video-LLMs) designed to enhance spatial-temporal modeling and audio understanding in video and audio-oriented tasks. Building upon its predecessor, VideoLLaMA 2 incorporates a tailor-made Spatial-Temporal Convolution (STC) connector, which effectively captures the intricate spatial and temporal dynamics of video data. Additionally, we integrate an Audio Branch into the model through joint training, thereby enriching the multimodal understanding capabilities of the model by seamlessly incorporating audio cues. Comprehensive evaluations on multiple-choice video question answering (MC-VQA), open-ended video question answering (OE-VQA), and video captioning (VC) tasks demonstrate that VideoLLaMA 2 consistently achieves competitive results among open-source models and even gets close to some proprietary models on several benchmarks. Furthermore, VideoLLaMA 2 exhibits reasonable improvements in audio-only and audio-video question-answering (AQA & OE-AVQA) benchmarks over existing models. All models are public to facilitate further research.
@@ 1 Introduction
In recent years, the field of artificial intelligence (AI) has achieved significant advancements (OpenAI, 2023a; Google, 2023; Anthropic, 2024), profoundly transforming industries and societal functions across the board. Models capable of image recognition (Bai et al., 2023; Dong et al., 2024; Chen et al., 2023d; Liu et al., 2024c; Chen et al., 2023b; Lin et al., 2023c; Young et al., 2024) and photorealistic image generation (Esser et al., 2024; Saharia et al., 2022; Ramesh et al., 2021) have approached near-human capabilities, catalyzing major breakthroughs in sectors such as medical imaging (Li et al., 2024b; Tu et al., 2024) and autonomous driving (Xu et al., 2023b; Jin et al., 2023a). Despite these successes, the domain of video understanding and generation (Kondratyuk et al., 2023; Menapace et al., 2024; Kondratyuk et al., 2023; Brooks et al., 2024) remains relatively nascent. Unlike static images, videos incorporate temporal dynamics and synchronous audio streams, significantly enriching the information content. This integration of continuous audio-visual data complicates extracting and interpreting meaningful patterns, amplifying data complexity and introducing unique computational challenges.
While Image Large Language Models (Image-LLMs) (Alayrac et al., 2022; Li et al., 2023a; Zhu et al., 2023b; Liu et al., 2024c; Ye et al., 2023; Bai et al., 2023; Chen et al., 2023d; Dong et al., 2024) processing static images have matured with impressive capabilities, Video Large Language Models (Video-LLMs) (Li et al., 2023b; Zhang et al., 2023; Maaz et al., 2023a; Lin et al., 2023a; Wang et al., 2024c; Liu et al., 2024a) lag notably behind due to inherent complexities. The primary challenge in video understanding lies in the temporal dynamics—recognizing visual patterns while understanding changes over time and correlating these with synchronous audio inputs. These temporal dynamics complicate the accurate prediction of future states and the understanding of complex scenarios, such as interactions among multiple entities or subtle environmental changes. Moreover, the integration of audio with visual data, essential for comprehensive understanding, remains underdeveloped in current models, limiting their effectiveness.
@@ 2.1 Architecture
With an established foundation in the visual domain, our exploration extends into the auditory realm to enhance the multimodal capabilities of VideoLLaMA. Initially, audio signals undergo a preprocessing step where they are transformed into fbank spectrograms with 128 frequency bins. To effectively harness these preprocessed audio signals, we integrate BEATs (Chen et al., 2023c), a cutting-edge audio encoder, known for its exceptional ability to capture detailed audio features and temporal dynamics. These features are then processed through a MLP block with two linear layers to align with the dimension of LLMs, therefore providing a more cohesive understanding of the video content when combined with the visual and acoustic modalities. By incorporating BEATs into VideoLLaMA, we address the challenge of synchronizing audio-visual data points. The encoder’s ability to capture temporal dynamics aligns with the STC Connector employed in the visual branch, ensuring a seamless integration of audio-visual features.
@@
Like its predecessor, VideoLLaMA 2 adopts the instruction-following large language models (LLMs) as its language decoder. We do not extensively search the optimal LLMs for VideoLLaMA 2 but use Mistral-Instruct (Jiang et al., 2023), Mixtral-Instruct (Jiang et al., 2024) and Qwen2-Instruct (Qwen, 2024) across all of the experiments. We leave the exploration of other popular LLMs, such as Gemma-IT (Gemma et al., 2024) and LLaMA3-Instruct (Dubey et al., 2024), for future work. We also include the results of VideoLLaMA 2 with Qwen2-Instruct (Qwen, 2024) as language decoder in Section 5.
@@
Spatial-Temporal Convolution Connector
The overall architecture and the workflow of our STC connector are depicted in Figure 2. We primarily follow three principles when designing the video-language connector: 1) Maintaining the spatial-temporal order in the output visual tokens; 2) Reducing the number of spatial-temporal tokens; 3) Alleviating information loss during spatial-temporal downsampling. ❶ The first principle guides us to avoid using resampler architecture (Li et al., 2023a; Alayrac et al., 2022) because the resampling operation does not guarantee the preservation of spatial-temporal order. This may cause suboptimal convergence because autoregressive models (i.e., LLM backbone) are highly dependent on consistent token order between training and inference. Therefore, we only consider the operations of convolution or pooling when building our connector. ❷ According to the second principle, we insert 3D downsample operator to compress spatial-temporal tokens. ❸ To complement the information loss caused by the spatial-temporal downsampling, we insert RegStage (a strong convolution block (Radosavovic et al., 2020)) before and after spatial-temporal downsampling. To empirically investigate the effectiveness of the designs above, we establish a quick but reasonable architectural search using the training data from Video-LLaVA (Lin et al., 2023a). We select three representative video understanding benchmarks: Egoschema (Mangalam et al., 2024), MV-Bench (Li et al., 2023c) and ActivityNet-QA (Yu et al., 2019) as the testbed111See Section 5.1 for benchmark details and evaluation protocols. and the comparison results are listed in Table 1. As can be seen, 3D convolution together with the ReStage block (i.e., the green line), which forms our STC connector, works the best in terms of the average performance. Another interesting finding is that almost all of the 3D downsampling designs perform better than the 2D ones on Egoschema, suggesting that the early fusion of frame-level features is beneficial for long video understanding.
@@ 3 Training
| Video-Text | InternVid-10M (Wang et al., 2023b) | 10M | 650K | 6.5% |
@@ 3.1.1 Pre-training
For the pre-training stage, we utilize a large-scale, weakly labeled, web-crawled dataset of image-text and video-text pairs, sourced from several publicly accessible databases. The video-text sources include Panda-70M (Chen et al., 2024c), VIDAL-10M (Zhu et al., 2023a), WebVid-10M (Bain et al., 2021), and InternVid-10M (Wang et al., 2023b), while the image-text sources include CC-3M (Changpinyo et al., 2021) and DCI (Urbanek et al., 2023).
@@ 3.1.2 Multi-task Fine-tuning
In the second stage of multi-task fine-tuning, we incorporate high-quality, fine-grained multi-modal annotations using both video-text and image-text data. As summarized in Table 3 , VideoLLaMA 2 is fine-tuned on four tasks simultaneously. For video captioning, we use training samples from VideoChat (Li et al., 2023b) and in-house collected data. For video classification and VQA tasks, we utilize a mixture of publicly available datasets, including Kinetics-710 (Kay et al., 2017), SthSthv2 (Goyal et al., 2017), NExtQA (Xiao et al., 2021), CLEVRER (Yi et al., 2019), EgoQA (Fan, 2019), Tgif (Li et al., 2016), WebVidQA (Yang et al., 2021), RealworldQA (X.AI, 2024), and Hm3d (Ramakrishnan et al., 2021). To enhance instruction-following capabilities, we collect and integrate data from multiple frontier Video-LLMs, such as Valley (Luo et al., 2023), VideoChatGPT (Maaz et al., 2023b), VideoChat (Li et al., 2023b), VTimeLLM (Huang et al., 2023), and VideoChat2 (Li et al., 2023c) to improve generalization. Additionally, we mix image captioning data from ShareGPT4V (Chen et al., 2023b) and image VQA and instruction-following data from LLaVA (Liu et al., 2023b, a) to maintain the capabilities in understanding static visual concepts. To further improve the instruction-following abilities, we incorporate pure text data from Magpie (Xu et al., 2024b) and ALLaVA (Xu et al., 2023a; Chen et al., 2024a).
@@ 3.2.2 Multi-task Fine-tuning
The second stage of audio-language training aims to enhance the versatility and applicability of our model through multi-task learning, involving a variety of datasets tailored to different audio processing tasks. ClothoAQA (Lipping et al., 2022b), with about 1,500 entries, is utilized for refining our model’s capabilities in question answering based on audio cues, with each sample enriched by six associated questions sourced through crowdsourcing. The instruction tuning phase also capitalizes on the continued use of WavCaps (Mei et al., 2023), alongside AudioCaps (Kim et al., 2019) and Clotho (Drossos et al., 2020), which together contribute about 453,000 audio samples for audio-text training. MusicCaps (Agostinelli et al., 2023), comprising roughly 5,400 entries, extends our model’s exposure to the musical domain, specializing in music-based captioning. For the purpose of sound event classification, we engage with VGGSound (Chen et al., 2020), which offers over 190,000 labeled audio clips of diverse acoustic events. UrbanSound8k (Salamon et al., 2014) and ESC50 (Piczak, 2015) provide additional layers of complexity with 8000+ and 2,000 urban and miscellaneous environmental sounds, respectively. TUT2017 (Mesaros et al., 2017), TUT2016 (Mesaros et al., 2016b), and VocalSound (Gong et al., 2022) enrich our dataset collection with more than 4,000, 1100, and 15,000 samples respectively, focusing on general sound events and human vocal sounds classification, thus broadening our model’s acoustic perceptual skills.
@@ 3.3 Audio-Video Joint Trainng
The third phase shifts our focus towards the integration of audio and visual modalities, aiming to harness and understand the interactions between these two crucial aspects of multimodal perception. This phase incorporates audio-visual datasets including AVQA (Yang et al., 2022) and AVQA-music (Li et al., 2022b), featuring around 57,000 and 35,000 entries, respectively. These datasets are specifically designed for audio-visual question answering, challenging our model to not only perceive but also interpret cross-modal content effectively. The AVSD (Alamri et al., 2019) dataset, consisting of approximately 80,000 pairs, is pivotal for developing and refining audio-visual dialogue systems. Furthermore, VGGSound (Chen et al., 2020) reappears with a substantial subset dedicated to audio-visual classification tasks, reinforcing the dataset’s utility and versatility in our multimodal training framework. Additionally, we incorporate the AVInstruct (Ye et al., 2024a) dataset, an audio-visual instruction dataset emphasizing co-learning on dynamic audio-visual pairs to address diverse AVQA tasks.
In addition to these audio-visual datasets, we also leverage several video-only datasets that enrich our system’s capacity to process and analyze diverse forms of input. For video-text pairs, our resources include SthSthv2 (Goyal et al., 2017) with 20K entries, EgoQA (Fan, 2019) featuring 7.8K, Tgif (Li et al., 2016) contributing 24.7K, Kinetics-710 (Kay et al., 2017) incorporating 19.5K, and VideoChat2 (Li et al., 2023c) with 9.5K. Other video-text datasets such as CLEVRER (2K) (Yi et al., 2019) and Valley (35.5K) (Luo et al., 2023) enhance our model’s robustness in understanding video content in conjunction with textual cues. Additionally, NExTQA (Xiao et al., 2021), though smaller with 1K pairs, serves as a valuable supplement. To further enhance the image-text interaction, we incorporate image-text datasets such as ShareGPT4V (Chen et al., 2023b) with a substantial 82K entries and LLaVA (Liu et al., 2023b, a) using 62K pairs, providing the necessary frameworks for visual comprehension alongside textual analysis.
Alongside these, we incorporate a broad range of audio-text pair datasets to bolster our model’s multimodal capabilities. Specifically, we include AudioCaps (Kim et al., 2019) with 49.3K entries, VocalSound (Gong et al., 2022) offering 15K, and UrbanSound8k (Salamon et al., 2014) with 8K, which supply extensive audio-text data for urban and vocal soundscapes. Additionally, we utilize Clotho (Drossos et al., 2020) with 19.3K entries, Wavcaps (Mei et al., 2023) with 100K, MusicCaps (Agostinelli et al., 2023) providing 5.4K, and niche datasets such as TUT2017 (Mesaros et al., 2017) and TUT2016 (Mesaros et al., 2016b) with 4.6K and 1.1K entries respectively, which focus on various environmental sounds. Furthermore, ClothoAQA (Lipping et al., 2022b) with 1.5K pairs adds to our resources for query-answering capabilities within audio-text contexts. These datasets collectively enable our model to engage with a rich variety of sound types and contexts, making it adept at handling complex audio-related tasks.
Finally, our text dataset, evol-instruct (Xu et al., 2023a), consisting of 23.2K examples, rounds out our multimodal approach, facilitating a comprehensive training environment that ensures our models can effectively follow instructions and interpret a wide range of multimodal inputs.
@@
In this stage, audio tracks are extracted from videos and cut to align the video clips. These audio clips are then truncated or padded to the same duration as the audio-language traning stage. For videos that lack an audio track, we fill the waveform with zeros to ensure uniformity across all data samples. During this training phase, we sample data for each batch in a 2:1:1 ratio of audio-visual data to visual data to audio data, ensuring that all audio-visual data is covered in one epoch. The video encoder remains frozen while we optimize the audio/video projector and the audio encoder, alongside the unfrozen LLM. The training objectives align with the instruction tuning stage, ensuring a coherent and effective progression in our multimodal training approach. By effectively leveraging the synchronized audio-visual data, VideoLLaMA 2 achieves a deeper understanding of multimodal content, thereby enhancing its performance across a spectrum of multimedia analysis tasks.
4 Implementation
Our VideoLLaMA2 is built upon LLaVA 1.5 library (Liu et al., 2023a)222https://github.com/haotian-liu/LLaVA. Across all VideoLLaMA 2 variants, we use clip-large-336333https://huggingface.co/openai/clip-vit-large-patch14-336 as the primary visual encoder, though siglip-so400m-384444https://huggingface.co/google/siglip-so400m-patch14-384 is preferred in later variants due to its superior performance. For audio encoding, we utilize Fine-tuned BEATs_iter3+(AS2M)(cpt2)555https://1drv.ms/u/s!AqeByhGUtINrgcpj8ujXH1YUtxooEg?e=E9Ncea. The language decoders are initialized with either Mistral-7B-Instruct666https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.2, Mixtral-8x7B-Instruct777https://huggingface.co/mistralai/Mixtral-8x7B-Instruct-v0.1, or, in some model variants, Qwen2-7B-Instruct888https://huggingface.co/Qwen/Qwen2-7B-Instruct and Qwen2-72B-Instruct999https://huggingface.co/Qwen/Qwen2-72B-Instruct. The audio branch training utilizes the video LLM initialized using the weights of the VideoLLaMA 2 (7B) model, which is trained on the two-stage video branch with 16 frames. We do not do any hyperparameter tuning during both pre-training and fine-tuning. Instead, we empirically set the global batch size and the learning rate as 1,024 and 1e-3 for pre-training and 2,048 and 2e-5 for fine-tuning. For the video-only training stage, VideoLLaMA 2 is pre-trained for just one epoch, followed by a fine-tuning process lasting up to three epochs. In the audio-only training, we also pre-train VideoLLaMA 2 for a single epoch, but limit fine-tuning to two epochs. For the audio-visual joint training stage, the pre-trained model undergoes fine-tuning for up to two epochs.